# ul. Pasteura 5, 02-093 Warsaw, Poland

### v1

@@ 2.1.2 Intercommutation probability
The value of the intercommutation probability can also be affected by the large extra dimensions. In highly symmetric scenarios like those in OCallaghan:2010rlo ; OCallaghan:2010mtk $P$ will be suppressed by the ratio of the extra-dimensional volume to the string volume, as the strings would now be able to “miss” each other in the extra dimensions. For the values of $\mathcal{V}$ considered here, this would lead to a negligibly small $P$, seriously impacting the ability to even form a cosmic superstring network in the first place. However, the scenario in the back of our minds considers generic ingredients on model-building such as branes and fluxes. As remarked in Jackson:2004zg , [^12] In a more detailed model for cosmic superstring evolution one may introduce introduce a time dependence on the effective volume parameter $w$ and on $g_{s}$, which would then determine the resulting time-dependence of the various interaction coefficients. Other discussions on effective volumes probed by cosmic strings appear in Avgoustidis:2007ju ; Avgoustidis:2012vc . the position of a string along the extra dimensions is a modulus like others, and will presumably be stabilised in a realistic compactification, reducing the effective length along which the strings will be able to propagate. Taking this into account, $P$ can be estimated as
$P\simeq 0.5g_{s}^{2}\left[\frac{6\pi}{\log\left(A\mathcal{V}\right)}\right]^{3}\,,$ (4)
@@ 2.2 Properties of the network
In this work, we will make the assumption that a population of cosmic superstrings is present after inflation, in sufficient numbers to be able to form a network. A typical production mechanism for superstrings is brane-antibrane annihilation, whose attractive potential is often advocated as a mechanism for inflation in String Theory Dvali:1998pa ; Burgess:2001fx ; Jones:2002cv ; Sarangi:2002yt ; Jones:2003da ; Pogosian:2003mz ; Kachru:2003sx (See also Burgess:2022nbx ; Cicoli:2024bwq for more recent developments). More generally, any scenario achieving a thermalized energy density of order one or more in string units will generically excite fundamental string oscillator modes, rendering large, heavy strings. Depending on the details of the compact space and whether there are open strings [^13] This is related to the stability or instability of the strings discussed in Copeland:2003bj . in the spectrum (i.e: on whether there are branes present where this energy density is reached), the thermodynamics is dominated by either a bath of open strings (rendering a Hagedorn phase, recently discussed in Frey:2023khe ; Frey:2024jqy ) or by an order one number of long, closed strings. [^14] This is sometimes called a Hagedorn phase transition in the literature. This is however different from what was considered in the seminal paper of Atick and Witten Atick:1988si as the latter requires a finite string coupling whilst the computations rendering a long string are performed in the $g_{s}\to 0$ limit. There is by now evidence (see e.g Barbon:2004dd ) that the Atick-Witten transition leads to black hole nucleation, so both pictures are related by the Horowitz-Polchinski correspondence Horowitz:1996nw ; Horowitz:1997jc between black holes and highly excited strings, stating that highly excited strings become black holes as $g_{s}$ gets large. In this article we work in the limit where the highly excited strings can be described as such. We refer to Abel:1999rq ; Frey:2023khe and Appendix A of Frey:2024jqy for further details on string thermodynamics, and simply point out that the latter case, with an order one number of highly excited strings per Hubble patch, that is of interest to us in this article. More recently, it has also been suggested that, in the presence of a time-dependent string tension, if a small number of strings were nucleated from the vacuum they would be able to percolate and form a network Conlon:2024uob (see also Brunelli:2025ems ).
@@
It has been known since the early days of the subject that, whenever a network of cosmic strings is formed, it will quickly reach an attractor known as the scaling regime, where the energy density of the strings constitutes a fixed fraction of the background Kibble:1984hp ; Bennett:1985qt ; Austin:1993rg ; Vilenkin:2000jqa . This is a self-similar configuration where correlation length of the network grows proportionally to the Hubble radius. Modulo some assumptions to be reviewed in this section, a similar configuration is reached in the case of varying tension, with the only difference that the relative energy density in strings will now evolve proportionally to the tension. In formulae,
$\Omega_{\rm{strings}}(t)\equiv\frac{\rho_{\rm{strings}}}{3H^{2}}=\Omega_{\rm{strings}}(t_{0})\frac{\mu(t)}{\mu(t_{0})},$ (5)
@@
for any given reference time $t_{0}$. The scaling regime can be described by a semi-analytical model known as the Velocity One Scale (VOS) model, introduced in Martins:1996jp ; Martins:2000cs . The VOS model was recently extended to case of varying tension in Revello:2024gwa , and here we give a brief overview of the results.
@@ 2.2.1 The VOS model (with a time dependent tension)
The VOS model Martins:1996jp ; Martins:2000cs can be obtained by coarse-graining of the equations of the classical equations of motion for the strings, supplemented by information obtained through numerical simulations. Denoting by $\left\{X^{\mu},\zeta_{a}\right\}$ spacetime and worldsheet coordinates respectively, the motion of a single string with a (space)time-dependent tension is described by
$S=-\int d^{2}\zeta\,\mu\left(X^{\rho}(\zeta)\right)\sqrt{-\rm{det}\left(\gamma_{ab}\right)}\quad\quad\quad\quad\gamma_{ab}=\frac{\partial X^{\mu}(\zeta)}{\partial\zeta^{a}}\frac{\partial X^{\nu}(\zeta)}{\partial\zeta^{b}}g_{\mu\nu}.$ (6)
The corresponding equations of motion are Emond:2021vts ; Conlon:2024uob
$\frac{1}{\sqrt{-\gamma}}\partial_{a}\left(\sqrt{-\gamma}\gamma^{ab}g_{\rho\sigma}X^{\sigma}_{\,,b}\right)-\frac{1}{2}\partial_{\rho}g_{\lambda\sigma}\gamma^{ab}X^{\lambda}_{\,,a}X^{\sigma}_{\,,b}=0.$ (7)
@@
[…]
@@
If the background is dominated by a fluid with $\rho\sim a^{-n}$ and the string tension scales as $t^{-q}$, $v$ and $L$ obey the system of equations Revello:2024gwa
$\begin{cases}2\frac{{\rm d}L}{{\rm d}t}&=2HL(1+v^{2})+\tilde{c}v+\frac{L}{\mu}\frac{{\rm d}\mu}{{\rm d}t}v^{2},\\ \frac{{\rm d}v}{{\rm d}t}&=(1-v^{2})\left[\frac{k(v)}{L}-\left(2H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}\right)v\right],\end{cases}$ (10)
@@
where $H$ is the Hubble ratio (in terms of $dt=a(t)d\tau$), $\tilde{c}$ the loop chopping efficiency (defined later in (13)) and $k(v)$ a phenomenologically motivated function known as the momentum parameter. For our numerical estimates, we will take $\tilde{c}=0.23$, derived from simulations in Martins:2000cs . An analytical expression for $k(v)$ was also suggested in Martins:2000cs , based on a combination of analytical arguments and a comparison to simulations:
$k(v)=\frac{2\sqrt{2}}{\pi}\left(1-v^{2}\right)\left(1+2\sqrt{2}v^{3}\right)\frac{1-8v^{6}}{1+8v^{6}}.$ (11)
@@
While the original analysis was restricted to the constant tension case, Revello:2024gwa provided evidence that the same formula should also describe strings with varying tension. Finally, the loop chopping efficiency $\tilde{c}$ is a phenomenological parameter describing energy loss into loops.
@@
We refer to Section 4 of Revello:2024gwa for a full analysis of the dynamical system, which in general can have multiple attractors. However, for reasonable initial conditions, the late time solutions will always approach a scaling solution characterized by $L=\xi t$ and
$\xi=\sqrt{\frac{n}{8}\frac{k(\bar{v})\left(k(\bar{v})+\tilde{c}\right)}{\left(1-\frac{qn}{4}\right)\left(1-\frac{2}{n}\right)}}\quad\quad\quad\quad\bar{v}=\sqrt{\frac{n}{2}\frac{k(\bar{v})}{k(\bar{v})+\tilde{c}}\frac{1-\frac{2}{n}}{1-\frac{qn}{4}}},$ (12)
@@
for a power-law cosmology $a\sim t^{2/n}$ with varying tension $\mu\sim t^{-q}$ (See Revello:2024gwa for more details). [^15] Notice the notation switch from $q$ in this paper to $m$ in Revello:2024gwa . In turn, these determine the number density of long strings through (8). This will be a fundamental quantity to compute to compute the GW spectrum in Section 4. For the rest of this work, we will assume such a scaling regime has been reached after a short transient.
@@
We are now in a position to discuss the effect of varying the intercommutation probability, which is the main difference between cosmic strings and superstrings at the network level. The idea is that the loop-chopping parameter is reduced by a power of the intercommutation probability, $\tilde{c}\to\tilde{c}P^{1-\beta}$. Considering $\beta\neq 0$ takes into account the fact that, because the intercommutation probability is reduced (and so the number of interactions), there is an increase in the small scale structure of the string. This increase leads to further contacts between the points of the string, an effect which competes against the reduction of the probability. The VOS predicts $\beta=0$ (since the increase of small scale structure generates a new scale that is not present in the model, see Avgoustidis:2005nv ), whilst simulations indicate $\beta=1/2$ in flat space Sakellariadou:2004wq and $\beta=2/3$ in an expanding spacetime Avgoustidis:2005nv . It has been shown Avelino:2012qy that weakly coupled strings reach a scaling regime with $\xi\sim\tilde{c}$, so for our present purposes we will take the effect of varying the intercommutation probability $P$ as taking $\tilde{c}\to P^{1-\beta}\tilde{c}$ and $\xi\to P^{1-\beta}\xi$, in agreement with previous studies Avelino:2012qy ; Sousa:2016ggw . Recent analyses in the context of PTA data take $\beta=1/2$ NANOGrav:2023hvm ; Figueroa:2023zhu ; Ellis:2023tsl and $\beta=2/3$ Avgoustidis:2025svu . In Sec. 4 we will study the GW spectrum analytically for all $\beta$, but in Sec. 5 we will take $\beta=1/2$. There is another effect that we will not be considering in this article, namely the existence of $D$-strings and the possibility of forming bound states with junctions, for which the VOS has been extended in Avgoustidis:2007aa ; Avgoustidis:2009ke ; Pourtsidou:2010gu . This effect only affects the low-frequency part of the spectrum Avgoustidis:2009ke ; Avgoustidis:2025svu and will not play a major role in our discussion, although an interesting future direction would be to study it in the context of varying tension.
@@ 2.2.2 The behavior of loops
When long strings intersect, they can form closed loops with a probability $P$. One can estimate the energy injected into the loop population as Kibble:1984hp
$\frac{d\rho}{dt}=\tilde{c}\bar{v}\frac{\rho}{L}\,$ (13)
which also provides a definition of the loop chopping parameter. Larger loops are the most relevant for GW emission Siemens:2006yp , and numerical simulations Blanco-Pillado:2013qja show that a fraction $\mathcal{F}_{\alpha}\simeq 0.1$ of the emitted loops have length $\mathcal{\ell}(t_{i})=\alpha t_{i}$, with $\alpha\simeq 0.1$.
It is not known what the appropriate value of $\alpha$ is in the case of cosmic superstrings. In particular, Ref. Sousa:2016ggw argues that a smaller intercommutation probability should affect the initial size of the loops. This size is related to the correlation length $L$ via $\ell=\alpha_{L}L$ (which defines $\alpha_{L}$). Since $L=\xi t$ and $\xi$ depends on the intercommutation probability, it follows that $\alpha=\alpha_{L}\xi$ inherits a $P$-dependence. [^16] Note that assuming $\alpha_{L}$ constant implies that $\alpha$ inherits a dependence in the equation of state of the background and the rate of varying tension through $\xi$, c.f. Eq. (12). In Sec. 4 we study the behaviour of the spectrum for arbitrary $\alpha$ but in Sec. 5 we will take $\alpha=0.1$. As discussed in Sousa:2016ggw and Sec. 4 explicitly, this results in a shift in the frequency of features in the GW spectrum. To take into account this effect, we define the parameter $\chi$ so that, for cosmic superstrings,
$\alpha\to\alpha P^{\chi}\,,$ (14)
[…]
@@
$\mathcal{F}_{\alpha}\frac{d\rho}{dt}=\mu\gamma\ell\frac{dn}{dt}$ (15)
@@
$\frac{dn}{dt}=\frac{\mathcal{F}_{\alpha}}{\gamma}\frac{\tilde{c}\bar{v}}{\alpha_{L}}\frac{1}{L^{4}}\equiv\frac{\mathcal{F}_{\alpha}}{\alpha}\frac{C_{\rm{eff}}}{t^{4}}\,,$ (16)
[…]
$\frac{dn}{dt}\to\frac{dn}{dt}P^{-2(1-\beta)-\chi}\,.$ (17)
Once they are produced, the loops will then start oscillating and emitting gravitational radiation Vilenkin:1981bx ; Vachaspati:1984gt ; Turok:1984cn ; Burden:1985md ; Olum:1999sg ; Moore:2001px . The evolution of their length is given by the combination of the effects due to the varying tension and the loss of energy due to gravitational radiation. The former is obtained by integration of the equations of motion for the loop Conlon:2024uob ; Revello:2024gwa and the latter is given by
$\frac{dE_{\rm{GW}}}{dt}=-\Gamma G\mu^{2}\,,$ (18)
where $\Gamma$ is a phenomenological parameter describing energy loss to GWs. From simulations, it can be estimated as $\Gamma\simeq 50$ Blanco-Pillado:2017oxo . This number might have to be revisited for superstrings, but we will be taking it as a benchmark in what follows. Combined, these effects imply the following evolution of equation for the length of the loops:
$\frac{{\rm d}\ell}{{\rm d}t}=-\ell\frac{1}{2\mu}\frac{d\mu}{dt}-\Gamma G\mu\,.$ (19)
As first noticed in Conlon:2024uob , a peculiarity of (19) is that it allows loops to grow faster than the scale factor if the tension decreases quickly enough. This can lead to novel phenomena, such as percolation, which we will however not consider in this work. [^18] See Section 4.4 of Revello:2024gwa for estimates of the percolation time in our setting. With a power-law dependence of the tension on time, $\mu(t)=\mu_{0}\left(\frac{t_{0}}{t}\right)^{q}$, and an initial condition $\ell(t_{i})=\alpha P^{\chi}t_{i}$, Eq.(19) is solved by
$\ell(t)=\left(\frac{t}{t_{i}}\right)^{q/2}\left(\alpha P^{\chi}-\frac{\Gamma G\mu_{0}}{\frac{3q}{2}-1}\left(\frac{t_{0}}{t_{i}}\right)^{q}\right)t_{i}+\frac{\Gamma G\mu_{0}}{\frac{3q}{2}-1}\left(\frac{t_{0}}{t}\right)^{q-1}t_{0}\,,$ (20)
@@
$\ell(t)=\left[\alpha P^{\chi}-\Gamma G\mu_{0}\left(\frac{t_{0}}{t_{i}}\right)^{2/3}\log\left(\frac{t}{t_{i}}\right)\right]\left(\frac{t}{t_{i}}\right)^{1/3}t_{i}\,$ (21)

### v2

@@ 2.1.2 Intercommutation probability
The value of the intercommutation probability can also be affected by the large extra dimensions. In highly symmetric scenarios like those in [60, 61] $P$ will be suppressed by the ratio of the extra-dimensional volume to the string volume, as the strings would now be able to “miss” each other in the extra dimensions. For the values of $\mathcal{V}$ considered here, this would lead to a negligibly small $P$, seriously impacting the ability to even form a cosmic superstring network in the first place. However, the scenario in the back of our minds considers generic ingredients on model-building such as branes and fluxes. As remarked in [49], [^12] In a more detailed model for cosmic superstring evolution one may introduce a time dependence on the effective volume parameter $w$ and on $g_{s}$, which would then determine the resulting time-dependence of the various interaction coefficients. Other discussions on effective volumes probed by cosmic strings appear in [62, 63]. the position of a string along the extra dimensions is a modulus like others, and will presumably be stabilised in a realistic compactification, reducing the effective length along which the strings will be able to propagate. Taking this into account, $P$ can be estimated as [49] (see Eq. (8.1))
$P\simeq 0.5g_{s}^{2}\left[\frac{6\pi}{\log\left(A\mathcal{V}\right)}\right]^{3}\,$, (2.4)
@@ 2.2 Properties of the network
In this work, we will make the assumption that a population of cosmic superstrings is present after inflation, in sufficient numbers to be able to form a network. A typical production mechanism for superstrings is brane-antibrane annihilation, whose attractive potential is often advocated as a mechanism for inflation in String Theory [64, 65, 66, 67, 68, 69, 70] (See also [71, 39] for more recent developments). More generally, any scenario achieving a thermalized energy density of order one or more in string units will generically excite fundamental string oscillator modes, rendering large, heavy strings. Depending on the details of the compact space and whether there are open strings [^13] This is related to the stability or instability of the strings discussed in [47]. in the spectrum (i.e: on whether there are branes present where this energy density is reached), the thermodynamics is dominated by either a bath of open strings (rendering a Hagedorn phase, recently discussed in [72, 73]) or by an order one number of long, closed strings. [^14] This is sometimes called a Hagedorn phase transition in the literature. This is however different from what was considered in the seminal paper of Atick and Witten [74] as the latter requires a finite string coupling whilst the computations rendering a long string are performed in the $g_{s}\to 0$ limit. There is by now evidence (see e.g [75]) that the Atick-Witten transition leads to black hole nucleation, so both pictures are related by the Horowitz-Polchinski correspondence [76, 77] between black holes and highly excited strings, stating that highly excited strings become black holes as $g_{s}$ gets large. In this article we work in the limit where the highly excited strings can be described as such. We refer to [78, 72] and Appendix A of [73] for further details on string thermodynamics, and simply point out that the latter case, with an order one number of highly excited strings per Hubble patch, that is of interest to us in this article. More recently, it has also been suggested that, in the presence of a time-dependent string tension, if a small number of strings were nucleated from the vacuum they would be able to percolate and form a network [79] (see also [80]).
@@
It has been known since the early days of the subject that, whenever a network of cosmic strings is formed, it will quickly reach an attractor known as the scaling regime, where the energy density of the strings constitutes a fixed fraction of the background [81, 82, 83, 84]. This is a self-similar configuration where correlation length of the network grows proportionally to the Hubble radius. Modulo some assumptions to be reviewed in this section, a similar configuration is reached in the case of varying tension, with the only difference that the relative energy density in strings will now evolve proportionally to the tension. In formulae,
$\Omega_{\rm{strings}}(t)\equiv\frac{\rho_{\rm{strings}}}{3H^{2}}=\Omega_{\rm{strings}}(t_{0})\frac{\mu(t)}{\mu(t_{0})},$ (2.5)
@@
for any given reference time $t_{0}$. The scaling regime can be described by a semi-analytical model known as the Velocity One Scale (VOS) model, introduced in [85, 86]. The VOS model was recently extended to case of varying tension in [43], and here we give a brief overview of the results.
@@ 2.2.1 The VOS model (with a time dependent tension)
The VOS model [85, 86] can be obtained by coarse-graining of the equations of the classical equations of motion for the strings, supplemented by information obtained through numerical simulations. Denoting by $\left\{X^{\mu},\zeta_{a}\right\}$ spacetime and worldsheet coordinates respectively, the motion of a single string with a (space)time-dependent tension is described by
$S=-\int d^{2}\zeta\,\mu\left(X^{\rho}(\zeta)\right)\sqrt{-\rm{det}\left(\gamma_{ab}\right)}\quad\quad\quad\quad\gamma_{ab}=\frac{\partial X^{\mu}(\zeta)}{\partial\zeta^{a}}\frac{\partial X^{\nu}(\zeta)}{\partial\zeta^{b}}g_{\mu\nu}.$ (2.6)
The corresponding equations of motion are [87, 79]
$\frac{1}{\sqrt{-\gamma}}\partial_{a}\left(\sqrt{-\gamma}\gamma^{ab}g_{\rho\sigma}X^{\sigma}_{\,,b}\right)-\frac{1}{2}\partial_{\rho}g_{\lambda\sigma}\gamma^{ab}X^{\lambda}_{\,,a}X^{\sigma}_{\,,b}=0.$ (2.7)
@@
$v^{2}=\frac{\int d\zeta^{1}\,\varepsilon\dot{x}^{2}}{\int d\zeta^{1}\,\varepsilon},\quad\quad\rm{where}\quad\quad\varepsilon\equiv\sqrt{\frac{x^{\prime 2}}{1-\dot{x}^{2}}}.$ (2.9)
@@
If the background is dominated by a fluid with $\rho\sim a^{-n}$ and the string tension scales as $t^{-q}$, $v$ and $L$ obey the system of equations [43]
$\begin{cases}2\frac{{\rm d}L}{{\rm d}t}&=2HL(1+v^{2})+\tilde{c}v+\frac{L}{\mu}\frac{{\rm d}\mu}{{\rm d}t}v^{2},\\ \frac{{\rm d}v}{{\rm d}t}&=(1-v^{2})\left[\frac{k(v)}{L}-\left(2H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}\right)v\right],\end{cases}$ (2.10)
@@
where $H$ is the Hubble ratio (in terms of $dt=a(t)d\tau$), $\tilde{c}$ the loop chopping efficiency (defined later in (2.13)) and $k(v)$ a phenomenologically motivated function known as the momentum parameter. For our numerical estimates, we will take $\tilde{c}=0.23$, derived from simulations in [86]. An analytical expression for $k(v)$ was also suggested in [86], based on a combination of analytical arguments and a comparison to simulations:
$k(v)=\frac{2\sqrt{2}}{\pi}\left(1-v^{2}\right)\left(1+2\sqrt{2}v^{3}\right)\frac{1-8v^{6}}{1+8v^{6}}.$ (2.11)
@@
While the original analysis was restricted to the constant tension case, [43] provided evidence that the same formula should also describe strings with varying tension. Finally, the loop chopping efficiency $\tilde{c}$ is a phenomenological parameter describing energy loss into loops.
@@
We refer to Section 4 of [43] for a full analysis of the dynamical system, which in general can have multiple attractors. However, for reasonable initial conditions, the late time solutions will always approach a scaling solution characterized by $L=\xi t$ and
$\xi=\sqrt{\frac{n}{8}\frac{k(\bar{v})\left(k(\bar{v})+\tilde{c}\right)}{\left(1-\frac{qn}{4}\right)\left(1-\frac{2}{n}\right)}}\quad\quad\quad\quad\bar{v}=\sqrt{\frac{n}{2}\frac{k(\bar{v})}{k(\bar{v})+\tilde{c}}\frac{1-\frac{2}{n}}{1-\frac{qn}{4}}},$ (2.12)
@@
for a power-law cosmology $a\sim t^{2/n}$ with varying tension $\mu\sim t^{-q}$ (See [43] for more details). [^15] Notice the notation switch from $q$ in this paper to $m$ in [43]. In turn, these determine the number density of long strings through (2.8). This will be a fundamental quantity to compute to compute the GW spectrum in Section 4. For the rest of this work, we will assume such a scaling regime has been reached after a short transient.
@@
We are now in a position to discuss the effect of varying the intercommutation probability, which is the main difference between cosmic strings and superstrings at the network level. The idea is that the loop-chopping parameter is reduced by a power of the intercommutation probability, $\tilde{c}\to\tilde{c}P^{1-\beta}$. Considering $\beta\neq 0$ takes into account the fact that, because the intercommutation probability is reduced (and so the number of interactions), there is an increase in the small scale structure of the string. This increase leads to further contacts between the points of the string, an effect which competes against the reduction of the probability. The VOS predicts $\beta=0$ (since the increase of small scale structure generates a new scale that is not present in the model, see [88]), whilst simulations indicate $\beta=1/2$ in flat space [89] and $\beta=2/3$ in an expanding spacetime [88]. It has been shown [90] that weakly coupled strings reach a scaling regime with $\xi\sim\tilde{c}$, so for our present purposes we will take the effect of varying the intercommutation probability $P$ as taking $\tilde{c}\to P^{1-\beta}\tilde{c}$ and $\xi\to P^{1-\beta}\xi$, in agreement with previous studies [90, 91]. Recent analyses in the context of PTA data take $\beta=1/2$ [15, 16, 7] and $\beta=2/3$ [19]. In Sec. 4 we will study the GW spectrum analytically for all $\beta$, but in Sec. 5 we will take $\beta=1/2$. There is another effect that we will not be considering in this article, namely the existence of $D$-strings and the possibility of forming bound states with junctions, for which the VOS has been extended in [92, 93, 94]. This effect only affects the low-frequency part of the spectrum [93, 19] and will not play a major role in our discussion, although an interesting future direction would be to study it in the context of varying tension.
@@ 2.2.2 The behavior of loops
When long strings intersect, they can form closed loops with a probability $P$. One can estimate the energy injected into the loop population as [81]
$\frac{d\rho}{dt}=\tilde{c}\bar{v}\frac{\rho}{L}\,$ (2.13)
which also provides a definition of the loop chopping parameter. Larger loops are the most relevant for GW emission [95], and numerical simulations [96] show that a fraction $\mathcal{F}_{\alpha}\simeq 0.1$ of the emitted loops have length $\mathcal{\ell}(t_{i})=\alpha t_{i}$, with $\alpha\simeq 0.1$.
It is not known what the appropriate value of $\alpha$ is in the case of cosmic superstrings. In particular, Ref. [91] argues that a smaller intercommutation probability should affect the initial size of the loops. This size is related to the correlation length $L$ via $\ell=\alpha_{L}L$ (which defines $\alpha_{L}$). Since $L=\xi t$ and $\xi$ depends on the intercommutation probability, it follows that $\alpha=\alpha_{L}\xi$ inherits a $P$-dependence. [^16] Note that assuming $\alpha_{L}$ constant implies that $\alpha$ inherits a dependence in the equation of state of the background and the rate of varying tension through $\xi$, c.f. Eq. (2.12). In Sec. 4 we study the behaviour of the spectrum for arbitrary $\alpha$ but in Sec. 5 we will take $\alpha=0.1$. As discussed in [91] and Sec. 4 explicitly, this results in a shift in the frequency of features in the GW spectrum. To take into account this effect, we define the parameter $\chi$ so that, for cosmic superstrings,
$\alpha\to\alpha P^{\chi}\,$, (2.14)
reproducing [91] for $\chi=1/3$. Recent studies [15, 16, 7] take $\chi=0$, while Ref. [19] considers $\chi\neq 0$. [^17] We cannot provide a value for $\chi$ that reproduces [19] because we work under the simplifying assumption that the scaling regime is achieved throughout cosmic history. Our parametrization is therefore useful in unifying and potentially comparing the approaches. In Sec. 4 we will discuss the dependence of the amplitude of the GW spectrum and transition frequencies in $\chi$, while in Sec. 5 we will take the benchmark of [7] and allow for different values of $\chi$.
@@
$\mathcal{F}_{\alpha}\frac{d\rho}{dt}=\mu\gamma\ell\frac{dn}{dt}$ (2.15)
@@
$\frac{dn}{dt}=\frac{\mathcal{F}_{\alpha}}{\gamma}\frac{\tilde{c}\bar{v}}{\alpha_{L}}\frac{1}{L^{4}}\equiv\frac{\mathcal{F}_{\alpha}}{\alpha}\frac{C_{\rm{eff}}}{t^{4}}\,$, (2.16)
where the last equality defines $C_{\rm{eff}}$. This number has been computed in [97] to be dependent on the equation of state of the Universe, with $C_{\rm{eff}}\in\{0.39,\,5.4,\,30.6\}$ during matter, radiation and kination domination, respectively. We will assume these values for the case where the tension varies in Sec. 5, and we note for later reference that the loop production rate scales with intercommutation probability like
$\frac{dn}{dt}\to\frac{dn}{dt}P^{-2(1-\beta)-\chi}\,.$ (2.17)
Once they are produced, the loops will then start oscillating and emitting gravitational radiation [98, 99, 100, 101, 102, 103]. The evolution of their length is given by the combination of the effects due to the varying tension and the loss of energy due to gravitational radiation. The former is obtained by integration of the equations of motion for the loop [79, 43] and the latter is given by
$\frac{dE_{\rm{GW}}}{dt}=-\Gamma G\mu^{2}\,$, (2.18)
where $\Gamma$ is a phenomenological parameter describing energy loss to GWs. From simulations, it can be estimated as $\Gamma\simeq 50$ [104]. This number might have to be revisited for superstrings. We will keep it arbitrary in our analytic estimates and choose $\Gamma=50$ for the plots. Combined, these effects imply the following evolution of equation for the length of the loops:
$\frac{{\rm d}\ell}{{\rm d}t}=-\ell\frac{1}{2\mu}\frac{d\mu}{dt}-\Gamma G\mu\,.$ (2.19)
As first noticed in [79], a peculiarity of (2.19) is that it allows loops to grow faster than the scale factor if the tension decreases quickly enough. This can lead to novel phenomena, such as percolation, which we will however not consider in this work. [^18] See Section 4.4 of [43] for estimates of the percolation time in our setting. With a power-law dependence of the tension on time, $\mu(t)=\mu_{0}\left(\frac{t_{0}}{t}\right)^{q}$, and an initial condition $\ell(t_{i})=\alpha P^{\chi}t_{i}$, Eq.(2.19) is solved by
$\ell(t)=\left(\frac{t}{t_{i}}\right)^{q/2}\left(\alpha P^{\chi}-\frac{\Gamma G\mu_{0}}{\frac{3q}{2}-1}\left(\frac{t_{0}}{t_{i}}\right)^{q}\right)t_{i}+\frac{\Gamma G\mu_{0}}{\frac{3q}{2}-1}\left(\frac{t_{0}}{t}\right)^{q-1}t_{0}\,$, (2.20)
@@
$\ell(t)=\left[\alpha P^{\chi}-\Gamma G\mu_{0}\left(\frac{t_{0}}{t_{i}}\right)^{2/3}\log\left(\frac{t}{t_{i}}\right)\right]\left(\frac{t}{t_{i}}\right)^{1/3}t_{i}\,$ (2.21)